the length of a rectangle is 1 cm longer than its width. if the diagonal of the rectangle is 4c, what are the demensions (l and w) of the rectangle?
Answer by venugopalramana(3286)

LET WIDTH =B
LENGTH =B+1
DIAGONAL =SQRT (L^2+B^2)=SQRT(B^2+(B+1)^2)=5...IS IT 4 OR 4C ?I AM TAKING IT AS 5
TO SHOW YOU THE METHOD.....SQUARING
B^2+B^2+2B+1=25
2B^2+2B-24=0
B^2+B-12=0
B^2+4B-3B-12=0
B(B+4)-3(B+4)=0
(B+4)(B-3)=0
B=3
L=4